OBSERVATION AND ANALYSIS OF IR AND RAMAN SPECTRA OF SOME SOLUTIONS AT HIGH PRES-SURES

R. AKHMEDJONOV, F. GANIEV, \underline{K.KHALILOV} and M. KULIEVA, Department of Physics, Samarkand State University, Samarkand, Uzbekistan 703004.

Simultaneous transitions in IR absorption spectra of binary and ternary solutions with CO2, CS2, CHCl3 and alcohols at high pressures and temperatures were measured. With the purpose of study of a degree of effect of pressure on a half-width and intensity in solutions of Raman spectra lines 606cm-1 of benzene, 261cm-1 of chloroform and their 90% solutions in cyclohexane were studied. In solutions the degree of reduction of a half-width of lines is little bit more than in pure liquids, in identical pressures range. We revealed the correlation of parameters of the simultaneous transition bands and Raman spectra.